OpenStudy (jiteshmeghwal9):

Let x, y, z, w satisfy \[\frac{x^2}{4^2-1^2}+\frac{y^2}{4^2-3^2}+\frac{z^2}{4^2-5^2}+\frac{w^2}{4^2-7^2}=1\]\[\frac{x^2}{6^2-1^2}+\frac{y^2}{6^2-3^2}+\frac{z^2}{6^2-5^2}+\frac{w^2}{6^2-7^2}=1\]\[\frac{x^2}{8^2-1^2}+\frac{y^2}{8^2-3^2}+\frac{z^2}{8^2-5^2}+\frac{w^2}{8^2-7^2}=1\]\[\frac{x^2}{10^2-1^2}+\frac{y^2}{10^2-3^2}+\frac{z^2}{10^2-5^2}+\frac{w^2}{10^2-7^2}=1\] Then the value of \(\Large{x^2+y^2+z^2+w^2}\) is

OpenStudy (jiteshmeghwal9):

@HolsterEmission @ParthKohli

OpenStudy (loser66):

132

OpenStudy (eliesaab):

Agree with 132

OpenStudy (eliesaab):

Change the variables for x1=x^2, y1=y^2,z1=z^2,w1=w^2, You will obtain 4 linear equations in 4 unknown. Solve and find the sum

OpenStudy (eliesaab):

There must be an easier way than the one I described above

OpenStudy (jiteshmeghwal9):

That's a long method which i won't try to use during examination.

OpenStudy (kainui):

I'd plug it into my TI-84 and invert the matrix to get the answer on my calculator in less than a minute if I was allowed.

OpenStudy (jiteshmeghwal9):

We are not given more than 2 minutes for solving questions in jee advanced examination

OpenStudy (eliesaab):

There is a unique set of scalars a,b,c,d to make the a* eqn1 + b* eqn2 + c*eqn3 +d*eqn4= x+y+z+w=a+b+c+d=132 This also requires solving 4 equations in 4 unknowns. @jjiteshmeghwal9 , if you know an easier way, please post it. Many people would love to see it

OpenStudy (jiteshmeghwal9):

U r right there is no shortcut method for this question but a less time consuming approach is by assuming 4^2 as t in equation 1 and then u will get a quadratic equation